16 a Find the y-intercept of the tangent to f (x) = x In x at the point where: i x = 1 it x = 2 iii x = 3. b Make a conjecture about the y-intercept of the tangent to f (x) = x In x at the point where = a, a > 0. c Prove your conjecture algebraically. 17 Find the axes intercepts of the tangent to y = x2 ex at x = 1. 18 Find the exact area of the shaded triangle. 



1
Expert's answer
2022-02-22T03:22:28-0500

16.

a.

"f'(x)=(x\ln x)'=\ln x+x(1/x)=\ln x+1"

The equation of the tangent line at "x=x_0"


"y=f'(x_0)x-f'(x_0)x_0+f(x_0)"

i. "x=1"


"f(1)=1\ln(1)=0"

"f'(1)=\ln (1)+1=1"

The equation of the tangent line at "x=1"

"y=x-1"

"x=0, y=-1"

"y" -intercept: "(0, -1)"


ii. "x=2"


"f(2)=2\ln(2)"

"f'(2)=\ln (2)+1"

The equation of the tangent line at "x=2"

"y=(\ln(2)+1)x-2\ln (2)-2+2\ln (2)"

"y=(\ln(2)+1)x-2"

"x=0, y=-2"

"y" -intercept: "(0, -2)"


iii. "x=3"


"f(3)=3\ln(3)"

"f'(3)=\ln (3)+1"

The equation of the tangent line at "x=3"

"y=(\ln(3)+1)x-3\ln (3)-3+3\ln (3)"

"y=(\ln(3)+1)x-3"

"x=0, y=-3"

"y" -intercept: "(0, -3)"


b.

We see that he "y"-intercept of the tangent to "f (x) = x \ln x" at the point where "x= a, a > 0" is "(0, -a)."


c. The equation of the tangent line at "x=a, a>0"


"y=f'(a)x-f'(a)a+f(a)"


Substitute


"y=f'(a)x-f'(a)a+f(a)"

"=(\ln(a)+1)x-a\ln (a)-a+a\ln(a)"

"=(\ln(a)+1)x-a"

When "x=0," "y=-a."

"y" -intercept: "(0, -a)"


17.


"f'=(x^2e^x)'=2xe^x+x^2e^x"

The equation of the tangent line at "x=x_0"


"y=f'(x_0)x-f'(x_0)x_0+f(x_0)"

"x=1"


"f(1)=e"

"f'(1)=3e"

The equation of the tangent line at "x=1"

"y=3ex-2e"

"x=0, y=-2e"

"y" -intercept: "(0, -2e)"


"y=0, 0=3ex-2e"

"x=2\/3"

"x" -intercept: "(2\/3, 0)"


18.


"f'=(3xe^{x/2})'=3e^{x/2}+\dfrac{3}{2}xe^{x/2}"

The equation of the tangent line at "x=x_0"

"y=f'(x_0)x-f'(x_0)x_0+f(x_0)"

"x=-1"

"f(-1)=-3e^{-1\/2}"

"f'(-1)=\dfrac{3}{2}e^{-1/2}"

The equation of the tangent line at "x=-1"

"y=\dfrac{3}{2}e^{-1/2}x-\dfrac{3}{2}e^{-1/2}"

"x=0, y=-\dfrac{3}{2}e^{-1/2}"

"y" -intercept: "(0, -\dfrac{3}{2}e^{-1/2})"

"y=0, 0=\dfrac{3}{2}e^{-1/2}x-\dfrac{3}{2}e^{-1/2}"

"x=1"

"x" -intercept: "(1, 0)"

"Area=\frac{1}{2}(1)(\dfrac{3}{2}e^{-1/2})=\dfrac{3}{4}e^{-1/2}({units}^2)"